OpenStudy (anonymous):

Everett is rolling a block with the numbers 1, 2, 3, 4, 5, and 6 on it. Finley is drawing one disk from a basket of three disks: one blue, one red, and one yellow. What is the probability that Everett and Finley end up with an even number and a blue disk? A 1/3 B 5/6 C 1/5 D 1/6 medals if right

OpenStudy (anonymous):

@texaschic101 @iGreen @Michele_Laino

OpenStudy (anonymous):

@candygurl000 @_jazmineee @nicoleee23

OpenStudy (michele_laino):

here the probability to get an even number is: favorable cases/ possible cases= 3/6=...?

OpenStudy (anonymous):

2

OpenStudy (michele_laino):

please 3/6=...?

OpenStudy (anonymous):

3/6=0.5

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, the probability, to get a blue disk, is: favorable cases/possible cases= 1/3

OpenStudy (anonymous):

So A?

OpenStudy (michele_laino):

so, the requested probability is given by the product of the two above probabilities, namely: (1/2)*(1/3)=...?

OpenStudy (michele_laino):

\[\large \frac{1}{2} \times \frac{1}{3} = ...?\]

OpenStudy (anonymous):

1/6 Would be my answer

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (anonymous):

Can you help me with a few more?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Nathan is flipping a coin three times. What is the probability that it will land on tails twice and heads once? https://app30.studyisland.com/pics/131712coin1.png

OpenStudy (michele_laino):

here, we have to understand what are the possible cases, and what are the favorable cases

OpenStudy (anonymous):

???? What is that?

OpenStudy (michele_laino):

the probability to get a tail is 1/2, and the probability to get a head is 1/2. Now the probability to get a tail at first flipping, is: 1/2 the probability to get another tail at second flipping is (1/2)*(1/2)=1/4 the probability to get a head at third flipping is: (1/2)*(1/2)*(1/2)=1/8

OpenStudy (michele_laino):

I have used your diagram

OpenStudy (michele_laino):

please, tell me when I may continue

OpenStudy (anonymous):

You can continue

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, there is more than one combination of two tails and one head, precisely those combinations are three

OpenStudy (anonymous):

Ok

OpenStudy (michele_laino):

so the requested probability is: \[\large 3 \times \frac{1}{8} = \frac{3}{8}\]

OpenStudy (anonymous):

Thank you please let me know if you don't want to do anymore

OpenStudy (michele_laino):

I'm ready!

OpenStudy (anonymous):

Directions: Select all the correct locations on the table. Spencer created a game where the player randomly draws a colored marble from a bag and rolls a standard six-sided die. In the bag, there are two red marbles, one blue marble, and two green marbles. One of the following tables correctly shows all of the possible outcomes for drawing a marble and rolling a die. Select each outcome in this table that represents drawing a green marble and rolling an odd number.

OpenStudy (anonymous):

Ther is the whole question

OpenStudy (anonymous):

z

OpenStudy (michele_laino):

the requested outcomes, are in the form (G, *) where * is an odd number

OpenStudy (anonymous):

I'm not understanding

OpenStudy (michele_laino):

we have to choose the right table, first

OpenStudy (anonymous):

Ok

OpenStudy (michele_laino):

after that we have to choose only the outcomes, like this: (G,*) where * stands for an odd number

OpenStudy (anonymous):

Ok

OpenStudy (michele_laino):

I'm choosing the right table...

OpenStudy (anonymous):

Ok

OpenStudy (michele_laino):

I think it is the table in the middle

OpenStudy (michele_laino):

so, being the requested outcomes: (G,1), (G,3), (G,5) how many of such outcomes, we can select from the table in the middle?

OpenStudy (anonymous):

Thats ok the extra practice timed out

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

do you need more help?

OpenStudy (anonymous):

Fill in the blank. The probability of a chance event is a number between 0 and ___ that expresses the likelihood of the event occurring.

OpenStudy (anonymous):

1. Fill in the blank. The probability of a chance event is a number between 0 and ___ that expresses the likelihood of the event occurring.

OpenStudy (anonymous):

Thois is the question

OpenStudy (anonymous):

This

OpenStudy (michele_laino):

hint: the probability can not be a number greter than 1

OpenStudy (michele_laino):

greater*

OpenStudy (anonymous):

So 1/2

OpenStudy (michele_laino):

1/2 is a possible value for a probability

OpenStudy (anonymous):

Answers 1/2 1 100 1/10

OpenStudy (michele_laino):

1/10 is a possible value for probability, whereas 100 is not a possible value for probability, since 100>1